OpenStudy (anonymous):

Solve and show the work. (3-n)/(2)=(3)/(4)

OpenStudy (anonymous):

(3-n)/(2)=(3)/(4) denominator are 2 and 4. LCM of them is 4 so, multiply both sides with 4 4(3-n)/2=(4)(3/4) 2(3-n)=3 expand it 6-2n=3 minus both sides with 6 in order to put x on one side and number at the other side 6-2n-6=3-6 -2n=-3 divide both sides by -2 (-2n)/(-2)=(-3)/(-2) n=3/2